OpenStudy (anonymous):

convert y=-4/3x-2 into standard form

OpenStudy (amistre64):

time it all by 3: 3y = -4x -6 ; add out the 4x 4x +3y = -6

OpenStudy (amistre64):

those steps arent in any particular order; you could have added out the -4/3x first: 4/3 x + y = -2 ; then times it all by 3 4x +3y = -6 and get the same answer